, The length of a seconds pendulum on the earth Where. G Is 9.8 m/s2 ls 1m. The length of a seconds pendulum on a planet where g = 4.9 m/s2 will be (1)1m (2)2m(3) 0.5 m (4) 0.25 m

Grade:11

3 Answers

Venky
7 years ago
g is directly  proportional to the length    so l1/l2=g1/g2      l1=1       g1 =9.8     g2 = 4.9     
substituting   l2  =      0.5        I think you got answer  if yes    aapprove yes
Rahul Verma
7 years ago
In seconds pendulum time period = 2 seconds
length of pendulum = 1m on earth
therefore putting in formula,
T=2\pi \sqrt{\frac{l_{1}}{G_{1}}} we get
G_{1}=\pi ^{2}l_{1}
similarly
G_{2}=\pi ^{2}l_{2} (on the other planet)
dividing both the eq we get
\frac{G_{1}=\pi ^{2}l_{1}}{G_{2}=\pi ^{2}l_{2}} 
\frac{G_{1}=\\l_{1}}{G_{2}=l_{2}}
\frac{9.8=1}{4.9=l_{2}}
from here we get l2 = 0.5m
 
 
somi teez
6 years ago
hi,
we know , t=2 pi (sqrt(L/g))
it means t is g is directly proportional to length of the pendulum
so,
g1/g2 =Ll1 / L2
SUBSTITUTING THE VALUES WE GET,
L2 = 0.5 m.....answer
approve if u agree